In a triangle, the measure of the first angle is twice the measure of the second angle. The measure of the third angle is 12 degrees more than the measure of the second angle. what is the measure of each angle?
Answer by rfer(16322)

2x+x+x+12=180
4x+12=180
4x=168
x=42 degrees
2x=84 degrees
x+12=54 degrees